# MN-TSG: Continuous Time Series Generation with Irregular Observations

### v1

@@ E.1 Datasets
The used medical ECG datasets are ECG200, ECG5000 (ECG5K), ECGFiveDays (ECGFD), and TwoLeadECG (TLECG), which can be downloaded from the link222www.timeseriesclassification.com. Their sequence lengths are 96, 140, 136, and 82. The number of classes is 2, 5, 2, and 2, respectively. The data sizes are 200, 5000, 884, and 1162, respectively.
E.2 Implementation details
@@
For the TSG backbone, we follow the standard diffusion model architecture. The denoising network adopts a U-Net with four down- and up-sampling stages. Each stage contains two residual blocks and one cross-attention block, where residual blocks use two 1D convolution layers, and cross-attention employs 1D convolutions for input/output projections with eight attention heads. A middle block, placed between the down- and up-sampling paths, consists of two residual blocks and one attention block. In addition, we include input and output layers, each implemented with a single 1D convolution. SiLU is used as the activation function throughout. We train the model for 600 epochs with a batch size of 256, a learning rate of . More details can be found in our code link333https://anonymous.4open.science/r/MNTSG-79C2.
Appendix F Additional experimental results
F.1 Further discussions on MOE weights
Regarding positivity of the generated weights. During training, all expert weights are positive. In practice, the diffusion model learns this distribution well, and we have not observed negative weights in the dataset. As shown in Table 6, the minimum expert weight across all samples and datasets remains strictly above zero.
Regarding the sum-to-one property of the generated weights. We do not enforce a hard constraint to force the generated weights to sum to 1. Instead, we allow the diffusion model to learn this property implicitly—a form of adaptive soft constraint. As reported in Table 7, the sum of the generated weights is extremely close to 1 across all datasets, indicating that the model naturally preserves this property without explicit enforcement. Moreover, Table 9 compares forecasting performance using (i) raw generated weights and (ii) post-normalized weights that sum to 1. The results show almost no difference between the two settings. This suggests that the learned joint distribution already produces weight vectors that are appropriate for each sample. Finally, Figure 8 and Figure 9 visualize the effect of normalizing the weights. The qualitative results are fully consistent with the above quantitative findings: whether or not the weights are normalized to sum to 1 has no noticeable impact on the generated trajectories.
Average weight of single expert. Table 8 reports the average weight assigned to each expert across all samples in different datasets. The weights are not sparse, and different datasets both tend to prefer different experts, demonstrating that the expert structure is indeed utilized and optimized in a meaningful manner rather than collapsing to a single expert.
| Sines | Stocks | Energy | MuJoCo | ||
| Global minimum | 0.1566 | 0.1553 | 0.1345 | 0.1393 |
| Sines | Stocks | Energy | MuJoCo | ||
| Mean | 1.0114 | 1.0046 | 1.0089 | 1.0085 | |
| Maximum | 1.0255 | 1.0305 | 1.0235 | 1.0163 | |
| Minimum | 0.9894 | 0.9766 | 0.9887 | 0.9726 | |
| Standard deviation | 0.0027 | 0.0044 | 0.0045 | 0.0068 |
| Expert 1 | Expert 2 | Expert 3 | Expert 4 | ||
| Sines | 0.31384414 | 0.2231202 | 0.26881662 | 0.19421977 | |
| Stocks | 0.19200034 | 0.21767974 | 0.32304975 | 0.2672697 | |
| Energy | 0.2413441 | 0.31066164 | 0.2065838 | 0.24141027 | |
| MuJoCo | 0.18345672 | 0.241069 | 0.2046078 | 0.370866 | |
| ECG200 | 0.21229412 | 0.44273788 | 0.20304906 | 0.14191894 | |
| ECG5K | 0.21999635 | 0.182932 | 0.3610456 | 0.2360259 | |
| ECGFD | 0.17836495 | 0.31938562 | 0.2739047 | 0.2283447 | |
| TLECG | 0.19264506 | 0.2577268 | 0.3210036 | 0.2286246 |
| 30% | 50% | 70% | |||||||||||
| MSE | Original | 0.026 | 0.002 | 0.013 | 0.029 | 0.026 | 0.002 | 0.013 | 0.029 | 0.028 | 0.002 | 0.013 | 0.029 |
| Normalize | 0.026 | 0.002 | 0.013 | 0.029 | 0.026 | 0.002 | 0.013 | 0.029 | 0.028 | 0.002 | 0.014 | 0.029 | |
| MAE | Original | 0.125 | 0.025 | 0.075 | 0.129 | 0.126 | 0.026 | 0.073 | 0.131 | 0.128 | 0.025 | 0.075 | 0.131 |
| Normalize | 0.125 | 0.025 | 0.075 | 0.129 | 0.126 | 0.026 | 0.073 | 0.131 | 0.128 | 0.025 | 0.075 | 0.131 | |
| 30% | 50% | 70% | ||||||||||||
F.2 Visualization Analysis of Continuous Generation
As shown in Figures 17, 18 and 19, from the visualization of continuous generation, it can be observed that the time series generated by our method generally conforms to the distribution of the original temporal patterns. Based on the dynamic functions learned during training, our method generates contextually relevant local fluctuation patterns at finer time intervals. This provides richer fine-grained information for downstream task decision-making, thereby contributing to improved accuracy in downstream prediction tasks.
This advantage is attributed to two main aspects. First, our proposed decoupled MOE NCDE enables NCDE to focus on learning multiple dynamic functions, adapting to complex and changing temporal patterns. This enhances its ability to handle irregular time series data and serve continuous generation tasks effectively. Second, by allowing the model to learn the joint distribution of mixture-of-experts dynamic functions and time series samples, it assigns new and appropriate expert weights to the generated time series samples, which are further used to weight the dynamic functions for continuous generation.
The visualization results demonstrate the effectiveness of the continuous time series generation method we proposed. In contrast, the continuous generation performances of GTGAN (Figures 20, 21 and 22) and KOVAE (Figures 23, 24 and 25) are relatively poor. The data they generate significantly deviates from the original temporal patterns, introducing substantial noise. This, in turn, leads to worse performance in downstream prediction tasks.
F.2.1 Hyperparameter sensitivity analysis on Discriminative score metric
F.2.2 Irregular time series generation
Results on each sequence length
In Table 1 of the main text, we reported the average results for time lengths of 12, 24, and 36. Here, we present the results of different methods for each length, as shown in Tables 11, 12, and 13. The conclusions are consistent with those in the main text. In the task of generating regular time series from irregular time series data, our method consistently outperforms the baseline models. This demonstrates that the proposed MOE-NCDE is effective. Through the mixture-of-experts dynamic functions and the decoupled training strategy, it enables NCDE to better capture dynamic temporal patterns, thereby better handling missing values and achieving the best results.
@@ F.2.2 Irregular time series generation
More qualitative evaluations
In the main text, we presented a qualitative study with 70% missing values in Figure 3. Here, we further demonstrate the performance with 30% and 50% missing values in Figure 27. The results show that the data distribution learned by our method is closer to the original data distribution than that of the current baseline models. This further illustrates the effectiveness of the proposed MOE-NCDE method.
@@
F.2.3 Continuous time series generation
Results of each sequence length on the downstream task
In Table 3 of the main text, we reported the average forecasting performance for time lengths of 12, 24, and 36. Here, we present the results of different methods for each length, as shown in Tables 14, 15, and 16. The conclusions drawn are consistent with those in the main text. Through refined continuous time series generation, our method has enabled downstream task models to achieve better forecasting performance. This is because the refined continuous generation provides more useful and richer information, which helps the models obtain better training effects.
More results for the analytical solutions of polynomial functions
As shown in Figure 28, we visualize the patterns of different curves in cubic polynomials. Since the polynomial coefficients are randomly sampled from normal distributions, respectively, there are differences in the patterns of different curves. The existing standard NCDE, which only uses a single dynamic function, has difficulty dealing with such dynamically changing temporal patterns. The dataset generation script can be found in the code link.
In the main text, we demonstrated the performance of our method in terms of analysis solution evaluation with 30% missing values. Here, we further show its performance with 50% and 70% missing values. As shown in Figure 29, after refined continuous generation, our method can still obtain good analysis solutions that conform to the true distribution. This further illustrates that our continuous generation method is effective and the generated finer-grained time values comply with the original data distribution. Additionally, we observe that continuous generation helps obtain better analysis solutions on coefficient , as shown in the fourth subplot of Figure 29 from left to right. By continuously generating refined curves (with a length twice that of the original), and then computing the analysis solutions on them, we can achieve more accurate solutions compared to normally generated samples. The enhanced accuracy is attributed to the richer information provided by the continuous generation process, which brings the distribution of the analysis solutions closer to the true distribution. This further illustrates the effectiveness of our continuous generation method.
We also present the comparison results of our method with KOVAE and GTGAN. As shown in Figures 30 and 31, since both KOVAE and GTGAN only use the standard NCDE, they have difficulty handling dynamically changing temporal patterns. The analysis solution distributions learned from the samples generated by them deviate from the true distribution. Our method, benefiting from the mixture-of-experts dynamic functions and decoupled design, can better learn different dynamic functions and thus better handle dynamically changing temporal patterns for more accurate interpolation.
Impact of control-path smoothness in NCDEs. Although the control path is smooth due to spline interpolation, the NCDE dynamics do not impose smoothness on the final generated trajectory. As shown in Eq. (1), the control path is smooth, but the evolution of is governed by the nonlinear neural network . Thus, even with a smooth control path, the latent trajectory can evolve in a highly non-smooth and expressive manner. Moreover, the mixture-of-experts component further facilitates the modeling of non-smooth and highly diverse temporal patterns. To validate this, we added visualizations of sequences generated solely by NCDE (Appendix Figures 14–16 in the revised manuscript). These results show:
For intrinsically smooth data (e.g., sine in Figure 16), NCDE produces smooth trajectories.
In summary, while the control path is smooth, the nonlinear MOE–NCDE dynamics allow the generated sequences to exhibit both smooth and abrupt temporal behaviors, depending entirely on the underlying temporal dynamics.
Discussions about generated patterns of MOE-NCDE. We provide additional visualizations in Appendix Figures 11 to 13. They show that the continuous trajectories generated by NCDE vary according to the learned dynamics and the MOE weights produced by the diffusion model. During continuous generation, NCDE exhibits diverse behaviors—including stable, fluctuating, and non-smooth patterns. The saw-tooth pattern is not representative of all cases but rather one instance of the broader range of dynamics the model can generate.
Discussion of how to select the expert number and its generalization. We recommend determining the number of experts based on two complementary principles:
Data-driven temporal-pattern analysis. We can perform clustering on dynamic features such as autocorrelation and volatility to estimate how many distinct temporal modes exist in a dataset. For our datasets (e.g., ECG with 2–5 disease classes, multivariate sine/stock datasets with periodic and aperiodic sub-patterns), four experts are sufficient to cover the dominant dynamic types without unnecessary model complexity.
Empirical validation. As shown in Figure 7 of the revised manuscript, we evaluate expert numbers ranging from 1 to 9. Performance consistently improves up to 4 experts, and then the benefit becomes negligible—adding more experts increases computation while offering no obvious benefit. Therefore, we select 4 experts as a balanced and generalizable default for TSG tasks.
Besides, Figure 7 further indicates that generalization performance is not highly sensitive to the expert number. While four experts provide the best overall results, using more experts does not significantly degrade performance, though it increases computational cost. Thus, the model’s generalization ability is stable with respect to this hyperparameter.

### v2

@@ D.1 Datasets
The used medical ECG datasets are ECG200, ECG5000 (ECG5K), ECGFiveDays (ECGFD), and TwoLeadECG (TLECG), which can be downloaded from the link111www.timeseriesclassification.com. Their sequence lengths are 96, 140, 136, and 82. The number of classes is 2, 5, 2, and 2, respectively. The data sizes are 200, 5000, 884, and 1162, respectively.
D.2 More implementation details
@@
When training MoE-NCDE, we use the Adam optimizer with a learning rate of , a batch size of 256, and a fixed hidden dimension of 64. For the TSG backbone of the standard diffusion model, the denoising network adopts a U-Net with four down- and up-sampling stages. Each stage contains two residual blocks and one cross-attention block, where residual blocks use two 1D convolution layers, and cross-attention employs 1D convolutions for input/output projections with eight attention heads. A middle block, placed between the down- and up-sampling paths, consists of two residual blocks and one attention block. In addition, we include input and output layers, each implemented with a single 1D convolution. SiLU is used as the activation function throughout. We train the model for 600 epochs with a batch size of 256, a learning rate of .
Except for the sequence length, all other dataset settings follow previous work [19]. More implementation details can be found in our code link222https://github.com/microsoft/TimeCraft/tree/main/Diff-MN. All model parameters used in the experiments are specified directly in the code.
Appendix E Supplementary experimental results
In Table 1 of the main text, we report the average performance of irregular-to-regular generation tasks across sequence lengths of 12, 24, and 36. Here, we additionally present results for each individual length. Overall, our method consistently outperforms others, in line with the conclusions from Table 1, and we therefore omit further discussion. We simply include them here for convenient reference.
E.1 Irregular to regular TSG on popular datasets with sequence length 12 (Table 6)
E.2 Irregular to regular TSG on popular datasets with sequence length 24 (Table 7)
E.3 Irregular to regular TSG on popular datasets with sequence length 36 (Table 8)
E.4 TSNE and PDF visualizations on irregular to regular TSG (Figure 9)
E.5 Evaluation of privacy protection capability (Table 9)
E.6 Visualizations about MoE-NCDE not producing overly smooth curves (Figure 13, 14 and 15)
E.7 Irregular to continuous TSG on popular datasets with sequence length 12 (Table 10)
E.8 Irregular to continuous TSG on popular datasets with sequence length 24 (Table 11)
E.9 Irregular to continuous TSG on popular datasets with sequence length 36 (Table 12)
E.10 Visualization of continuous TSG on the Stock dataset (Figure 10 for part 1, Figure 11 for part 2, and Figure 12 for part 3)
E.11 Visualization of cubic polynomial curves (Figure 16)
E.12 Cubic polynomial solutions of Diff-MN with 50% and 70% missing values (Figure 17)
E.13 Cubic polynomial solutions of Diff-MN VS. KOVAE with 30%, 50% and 70% missing values (Figure 19)
E.14 Cubic polynomial solutions of Diff-MN VS. GT-GAN with 30%, 50% and 70% missing values (Figure 20)
E.15 Visualization of temporal dynamics learned by different expert (Figure 18)
E.16 Average expert weights learned by MoE-NCDE (Figure 18)
E.17 The impact of the number of experts on the DS metric (Figure 8)
@@
| 30% | 50% | 70% | ||||||||||||
@@
| Datasets | Expert 1 | Expert 2 | Expert 3 | Expert 4 | |
| Sines | 0.31384414 | 0.2231202 | 0.26881662 | 0.19421977 | |
| Stocks | 0.19200034 | 0.21767974 | 0.32304975 | 0.2672697 | |
| Energy | 0.2413441 | 0.31066164 | 0.2065838 | 0.24141027 | |
| MuJoCo | 0.18345672 | 0.241069 | 0.2046078 | 0.370866 | |
| ECG200 | 0.21229412 | 0.44273788 | 0.20304906 | 0.14191894 | |
| ECG5K | 0.21999635 | 0.182932 | 0.3610456 | 0.2360259 | |
| ECGFD | 0.17836495 | 0.31938562 | 0.2739047 | 0.2283447 | |
| TLECG | 0.19264506 | 0.2577268 | 0.3210036 | 0.2286246 |